# Contextual Stochastic Optimization for Order Fulfillment Optimization

### v2

@@ 6.7.1 Flexibility of C-RO-B
Figures 7 and 8 summarize how these parameters affect solution quality under both C‑RO‑B‑CLF and C‑RO‑B‑QRF. As the prediction interval coverage increases from 70% to 90%. In Figure 7(a), increasing coverage from 70% to 90% reduces the average objective to about 0.84 (C‑RO‑B‑CLF) and 0.86 (C‑RO‑B‑QRF), an approximately 4% improvement relative to 70% coverage. This is consistent with the concurrent decline in cumulative lateness shown in Figure 7(c). However, pushing coverage beyond 95% reverses this trend: at 99% coverage the average objective rises sharply due to the cost of extreme conservatism and, for C‑RO‑B‑CLF, degraded prediction accuracy (reflected in a spike in lateness). By contrast, C‑RO‑B‑QRF continues to reduce lateness at 99%, illustrating its superior tail modeling. In terms of worst‐case performance (Figure 7(b)), both methods improve with coverage, but C‑RO‑B‑CLF plateaus after 95%, whereas C‑RO‑B‑QRF continues to improve at 99%.
@@ 6.9 Perturbation Analysis under Varying Simulation Environments
To assess the robustness of the proposed CSO methods, this section evaluates performance under varying simulation environments by introducing systematic perturbations to the simulated delivery time deviation distributions. These perturbations represent moderate to adversarial divergence from the test distribution, providing insight into the methods’ stability under distributional shifts.
Two types of perturbations are considered: random uniform noise and delay-specific bias applied to the simulation probability distributions.
@@ 6.10 Computational Scalability of the CSO Methods
Since single-item orders account for the majority of order volume, multi-item orders (what the methods are designed for) can be routed to a dedicated sourcing engine without disrupting standard operations. Under normal load, any of the three methods may be applied; under peak loads—such as holiday surges—C‑RO‑B is recommended to preserve fast solve times while maintaining high solution quality. Concurrent batches of orders can be processed in parallel when computational resources permit.
@@ 6.11 Managerial and Operational Implications
The proposed fulfillment framework offers actionable guidance for practitioners seeking to align sourcing strategies with demand variability, service-level requirements, and computational constraints. Depending on the operating context, different methods are preferable. When demand is stable and the goal is to minimize average fulfillment cost, C-SAA-CLF provides the most cost-effective solution. Under moderate runtime constraints, C-RO-B-CLF offers a balanced trade-off between efficiency and robustness. In contrast, during high-stakes periods such as holiday promotions or flash sales—where the cost of service-level violations is high—C-RO-D-CLF and C-RO-B-QRF (with 99% coverage) are better suited to mitigate tail risks and ensure reliability. This scenario-driven approach allows e-commerce platforms to dynamically adapt fulfillment strategies to shifting conditions while leveraging a unified optimization infrastructure. As retailers increasingly face variable demand patterns and rising customer expectations, the ability to flexibly match fulfillment strategy to context becomes a key operational advantage.
@@ 7 Conclusion
This paper introduces a generic contextual stochastic optimization (CSO) framework for data‑driven decision making when only observational data—hence missing counterfactual costs—are available. The framework is deliberately modular. First, any algorithm capable of returning a full predictive distribution can serve as the contextual distribution oracle. In the empirical study, calibrated multi‑class classifiers and tree‑based quantile‑regression forests were selected to capture the discrete, ordered nature of delivery‑time deviations. Second, the oracle feeds directly into two tractable optimization engines: (i) a Contextual Sample‑Average Approximation (C‑SAA) for risk‑neutral objectives and (ii) a Contextual Robust Optimization (C‑RO) formulation for risk‑averse objectives. Because the learning and optimization layers communicate only through the oracle, the template can be readily translated to other operations‐management settings.
@@
  • Multi‑period, inventory‑aware CSO: Extend the framework to jointly optimize sourcing, shipping, and stocking over rolling horizons, capturing inventory replenishment spillovers and elevate the model from a myopic to a long-term strategic planning tool.

### v3

@@ 6.7.1 Flexibility of C-RO-B
Figures 8 and 9 summarize how these parameters affect solution quality under both C‑RO‑B‑CLF and C‑RO‑B‑QRF. As the prediction interval coverage increases from 70% to 90%. In Figure 8(a), increasing coverage from 70% to 90% reduces the average objective to about 0.84 (C‑RO‑B‑CLF) and 0.86 (C‑RO‑B‑QRF), an approximately 4% improvement relative to 70% coverage. This is consistent with the concurrent decline in cumulative lateness shown in Figure 8(c). However, pushing coverage beyond 95% reverses this trend: at 99% coverage the average objective rises sharply due to the cost of extreme conservatism and, for C‑RO‑B‑CLF, degraded prediction accuracy (reflected in a spike in lateness). By contrast, C‑RO‑B‑QRF continues to reduce lateness at 99%, illustrating its superior tail modeling. In terms of worst‐case performance (Figure 8(b)), both methods improve with coverage, but C‑RO‑B‑CLF plateaus after 95%, whereas C‑RO‑B‑QRF continues to improve at 99%.
@@ 6.9 Perturbation Analysis under Varying Simulation Environments
A critical challenge in any simulation-based evaluation is that the simulator itself is an imperfect model of the real-world data-generating process. This is particularly true in this application, where the purely observational nature of the historical data, and the resulting lack of counterfactual outcomes, make it impossible to construct a perfect, clairvoyant simulator. Therefore, to rigorously assess the stability of the proposed solutions, this section performs a stress test designed to probe the framework’s performance under a misspecified ground truth. This analysis is conducted by introducing systematic perturbations to the simulated delivery time deviation distributions, representing moderate to adversarial divergences from the test data. Specifically, two types of perturbations are considered: random uniform noise and a delay-specific bias applied to the simulation probability distributions. Together, these tests provide critical insight into the methods’ stability under real-world distributional shifts and their robustness to the inevitable simulator-reality mismatch.
@@ 6.10 Computational Scalability of the CSO Methods
For completeness, a comparison of computation times against the baseline models is provided (see Appendix K). The Greedy and PTO methods are computationally cheaper, solving most instances in fractions of a second. However, the proposed CSO methods, particularly the highly-efficient C-RO-B, remain well within practical time limits for real-world deployment (Figure 13(a)). The substantial improvements in solution quality (Table 3) justify this modest increase in computational overhead, representing a favorable trade-off between performance and speed.
Since single-item orders account for the majority of order volume, multi-item orders (what the methods are designed for) can be routed to a dedicated fulfillment optimization engine without disrupting standard operations. Under normal load, any of the three methods may be applied; under peak loads—such as holiday surges—C‑RO‑B is recommended to preserve fast solve times while maintaining high solution quality. Concurrent batches of orders can be processed in parallel when computational resources permit.
@@ 6.11 Managerial and Operational Implications
The proposed fulfillment framework offers actionable guidance for practitioners seeking to align fulfillment strategies with demand variability, service-level requirements, and computational constraints. Depending on the operating context, different methods are preferable. When demand is stable and the goal is to minimize average fulfillment cost, C-SAA-CLF provides the most cost-effective solution. Under moderate runtime constraints, C-RO-B-CLF offers a balanced trade-off between efficiency and robustness. In contrast, during high-stakes periods such as holiday promotions or flash sales—where the cost of service-level violations is high—C-RO-D-CLF and C-RO-B-QRF (with 99% coverage) are better suited to mitigate tail risks and ensure reliability. This scenario-driven approach allows e-commerce platforms to dynamically adapt fulfillment strategies to shifting conditions while leveraging a unified optimization infrastructure. As retailers increasingly face variable demand patterns and rising customer expectations, the ability to flexibly match fulfillment strategy to context becomes a key operational advantage.
@@ 7 Conclusion
This paper introduces a generic contextual stochastic optimization (CSO) framework for data‑driven decision making when only observational data, hence missing counterfactual costs, are available. The framework is deliberately modular. First, any algorithm capable of returning a full predictive distribution can serve as the contextual distribution oracle. In the empirical study, calibrated multi‑class classifiers and tree‑based quantile‑regression forests were selected to capture the discrete, ordered nature of delivery‑time deviations. Second, the oracle feeds directly into two tractable optimization engines: (i) a Contextual Sample‑Average Approximation (C‑SAA) for risk‑neutral objectives and (ii) a Contextual Robust Optimization (C‑RO) formulation for risk‑averse objectives. Because the learning and optimization layers communicate only through the oracle, the template can be readily translated to other operations‐management settings.
@@
  • Multi‑period, inventory‑aware CSO: Extend the framework to jointly optimize fulfillment, shipping, and stocking over rolling horizons, capturing inventory replenishment spillovers and elevate the model from a myopic to a long-term strategic planning tool.
  • Generalization to broader problem classes: Extending the framework to broader settings like network flow optimization is a key direction for future work. The sampling-based C-SAA and C-RO-D approaches can be directly generalized. However, the C-RO-B formulation’s tractability relies on the specific problem structure, and adapting it to more complex constraints would require new reformulations.
{APPENDIX}
8 MILP Model Notations
Table 4 presents the notations used in the MILP optimization model.
Table 4: Optimization Model Nomenclature
Sets
$\mathcal{I}$ Set of SKUs, indexed by $i$.
$\mathcal{K}$ Set of carriers, indexed by $k$.
$\mathcal{L}$ Set of locations, indexed by $\ell$.
Parameters
$\operatorname{inv}_{i\ell}$ Inventory level of SKU $i$ at location $\ell$ when the order is placed.
$\operatorname{cap}_{\ell}$ Available capacity at location $\ell$ when the order is placed.
$e_{ik\ell}$ 1 if (carrier $k$, location $\ell$) is an eligible candidate location-carrier pair for SKU $i$ in the order and 0 otherwise.
$c_{ik\ell}^{ship}$ Per-unit shipping cost when using location $\ell$ and carrier $k$ to ship SKU $i$.
$c_{\ell}^{fixed}$ Per-unit fixed sourcing (non-parcel) cost when using location $l$.
$\beta_{k}$ Discount factor for shipping multiple units using carrier $k$.
$q_{i}$ Quantity of SKU $i$ in the order.
$d_{k\ell}$ Delivery time deviation when using location $\ell$ and carrier $k$.
$\gamma^{+}$ Constant that converts late delivery penalty into a per-unit cost.
$\gamma^{-}$ Constant that converts early delivery penalty into a per-unit cost.
NN A large number = $\sum_{i\in\mathcal{I}}q_{i}+2$.
$M_{k\ell}$ A large number = $\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}q_{i}$.
Decision Variables
$z_{ik\ell}$ Integer decision variable which measures the quantity of SKU $i$ in the order sourced from location $\ell$ and shipped by carrier $k$.
$y_{k\ell}$ Binary decision variable which equals 1 if more than one unit is sourced from location $\ell$ and carrier $k$ within the order, and equals 0 otherwise.
$w_{k\ell}$ Continuous decision variable which measures the cost to be discounted in the order using location-carrier pair $(k,\ell)$.
9 Data Preprocessing
To optimize the dataset for machine learning and optimization models, several preprocessing steps were implemented to clean the data. Duplicate entries and rows with missing values were removed. Geographic coordinates were generated from the zip codes of the fulfillment centers and customer locations, with any invalid zip codes discarded in the process. The geographic distance between customer and fulfillment center locations was approximated using the Haversine formula. Additionally, the dataset was filtered to include only carriers with sufficient order history for training prediction models, resulting in a subset of carriers.
10 Descriptive Analytics and Insights
The company receives tens of thousands of home delivery online orders daily, with an average ranging from 30,000 to 70,000 orders. During peak hours, 500 to 2,000 orders may be received within a 15-minute time window. Direct information regarding which carriers are available to ship specific SKUs from particular fulfillment centers is not available in the dataset. However, heuristic rules can be derived through analyses of historical orders. The customer’s location and the dimensions/weight of the SKU(s) in an order are critical factors in determining carrier suitability. Some carriers show a more scattered distribution of customer locations, indicating limited coverage in certain states of the U.S.. Additionally, coverage is dependent on the location of the fulfillment center from which the order is shipped, as some carriers can only serve a limited range around that center. In contrast, other carriers have served locations spanning nearly the entire US. Furthermore, the upper bounds of SKU length, height, and weight in the historical data reflect the package size and weight limits of different carriers.
Figure 14: Delivery Deviation Distribution of Orders Shipped by Three Example Carriers.
The distribution of deviations varies significantly across carriers, as illustrated in Figure 14. For the first two carriers, the majority of orders are delivered on time. The first carrier shows nearly all on-time deliveries, with the remaining orders arriving late. The second carrier has approximately half of the orders delivered on time, with slightly more early deliveries than late ones. In contrast, orders shipped by the last carrier predominantly arrive one day early, and most deliveries arrive ahead of schedule. Given these pronounced differences, separate prediction models were developed for each carrier.
11 Feature Importance
To identify the most important features for predicting deviations, a feature importance analysis was conducted using a tree-based regression model. Two methods were examined: Mean Decrease in Impurity (MDI) and Shapley values (SHAP). MDI measures the importance of a feature by calculating the total reduction in impurity (e.g., Gini impurity or entropy) it brings when used in decision trees (Breiman 2001). SHAP values, on the other hand, provide a unified measure of feature importance based on cooperative game theory, attributing to each feature its contribution to the prediction by considering all possible feature combinations (Lundberg et al. 2020).
Figure 15 highlights the top 10 most important features evaluated, as evaluated using MDI and SHAP, for an example carrier under a tree-based regression model. For this carrier, the planned lead time (promised delivery time) emerges as the most significant factor. Additionally, order information such as the release time and SKU type, fulfillment center characteristics like on-hand capacity, and carrier-specific attributes such as promised time-in-transit and carrier zone also play important roles.
(a) MDI Feature Importance.
(b) SHAP Value Feature Importance.
Figure 15: Top 10 Mean Regression Features for Delivery Deviation Measured by Two Feature Importance Metrics For an Example Carrier.
12 Linearize the Objective Function of the Order Fulfillment Problem
The nonlinear objective function $g(\mathbf{z},\mathbf{d})$ can be linearlized using auxiliary variables. Let $\mathbf{y}=(y_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}\in\{0,1\}^{(|\mathcal{K}||\mathcal{L}|)}$ be indicator variables, where $y_{k\ell}=1$ indicates more than one unit is sourced from location $\ell$ and shipped by carrier $k$ within the order. Additionally, let $\mathbf{w}=(w_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}\in\mathbb{R}^{(|\mathcal{K}||\mathcal{L}|)+}$ be continuous variables, where $w_{k\ell}$ represents the shipping costs to be discounted in the order using carrier-location pair ($k,\ell$).
The linear objective function and the additional constraints for linearization are defined as follows:
$g(\mathbf{z},\mathbf{d})=$ $(\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}(c_{\ell}^{fixed}+c_{ik\ell}^{ship}+\gamma^{+}d_{k\ell}^{+}+\gamma^{-}d_{k\ell}^{-})z_{ik\ell}-\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\beta_{k}w_{k\ell})$ (24)
$\begin{aligned} \text{s.t.}\quad&\sum_{i\in\mathcal{I}}z_{ik\ell}\geq 2-N(1-y_{k\ell})\\ &\sum_{i\in\mathcal{I}}z_{ik\ell}\leq 1+Ny_{k\ell}\\ &w_{k\ell}\leq M_{k\ell}y_{k\ell}\\ &w_{k\ell}\leq\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}z_{ik\ell}+(1-y_{k\ell})M_{k\ell}\\ &w_{k\ell}\geq\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}z_{ik\ell}-(1-y_{k\ell})M_{k\ell}\\ &y_{k\ell}\in\{0,1\}\\ &w_{k\ell}\geq 0\\ \end{aligned}\quad,\forall{k\in\mathcal{K}},{\ell\in\mathcal{L}},\\$ (25)
where $N$ and $M_{k\ell}$ are large constants for the big-M constraints, with their strictest values being $\sum_{i\in\mathcal{I}}q_{i}+2$ and $\sum_{i\in\mathcal{I}}c_{ik\ell}^{ship}q_{i}$, respectively.
13 Applying the C-RO-B to CSOFP
[…]
Substituting $\tilde{\mathbf{c}}$ leads to:
$\min_{\mathbf{z}\in\mathcal{Z}}\max_{\tilde{\mathbf{c}}\in\mathcal{U}_{b}(\mathbf{s})}$ $(\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\tilde{c}_{ik\ell}z_{ik\ell}-\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\beta_{k}w_{k\ell}).$ (26)
Given that the prediction interval for each $\tilde{d}_{k\ell}$ is $\underline{d}_{k\ell},\underline{d}_{k\ell}+\hat{d}_{k\ell}]$, where $\underline{d}_{k\ell}$ and ($\underline{d}_{k\ell}+\hat{d}_{k\ell}$) are the estimated lower and upper quantiles of the deviation for the carrier-location pair $(k,\ell)$, the prediction interval for $\tilde{\mathbf{c}}$ can be expressed as:
[…]
Let $B\in[0,(|\mathcal{K}||\mathcal{L}|)]$ be the uncertainty budget. The contextual budgeted interval uncertainty set $\mathcal{U}_{b}(\mathbf{s})$ is expressed as follows:
$\mathcal{U}_{b}(\mathbf{s},B)=\{\bm{\xi}\in\mathbb{R}^{(|\mathcal{I}||\mathcal{K}||\mathcal{L}|)}:\xi_{ik\ell}=\underline{\xi}_{ik\ell}+\hat{\xi}_{ik\ell}\delta_{k\ell},\sum_{k\in\mathcal{K},\ell\in\mathcal{L}}\delta_{k\ell}\leq B,\bm{\delta}\in[0,1]^{(|\mathcal{K}||\mathcal{L}|)}\}.$ (28)
Then, with fixed $\mathbf{z}$, the inner maximization problem of Problem (26) becomes:
$\max\quad$ $\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\hat{\xi}_{ik\ell}z_{ik\ell}\delta_{k\ell}$ (29)
s.t. $\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\delta_{k\ell}\leq B$ (30)
$\delta_{k\ell}\in[0,1],\quad\forall k\in\mathcal{K},\ell\in\mathcal{L}.$ (31)
The robust counterpart of Problem (26) is formulated as follows:
$\min_{\mathbf{z}\in\mathcal{Z}}$ $(\sum_{i\in\mathcal{I}}\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\underline{\xi}_{ik\ell}z_{ik\ell}-\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\beta_{k}w_{k\ell}+\pi B+\sum_{k\in\mathcal{K}}\sum_{\ell\in\mathcal{L}}\lambda_{k\ell})$ (32)
s.t. $\pi+\lambda_{k\ell}\geq\hat{\xi}_{ik\ell}z_{ik\ell},\quad\forall i\in\mathcal{I},k\in\mathcal{K},\ell\in\mathcal{L}$
$\pi\geq 0$
$\lambda_{k\ell}\geq 0,\quad\forall k\in\mathcal{K},\ell\in\mathcal{L},$
where $\pi$ and $(\lambda_{k\ell})_{k\in\mathcal{K},\ell\in\mathcal{L}}$ are the dual variables of Constraints (30) and (31), respectively.
14 Probabilistic Multi-Class Classification Using Ensemble Learning
RF leverages multiple classification trees trained on different random subsets of the data, making predictions by averaging the results (Breiman 2001). Let $\mathcal{R}_{1}^{(t)},\ldots,\mathcal{R}_{R}^{(t)}$ correspond to the partitioned regions of the $t$-th tree. Given covariates $\mathbf{s}$, the predicted probability for class $c$ in an RF classifier is:
$\hat{p}_{c}(\mathbf{s})=\frac{1}{T}\sum_{t=1}^{T}\frac{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}^{(t)}\}\cdot\mathbf{1}\{d=\xi_{c}\}}{\sum_{r=1}^{R}\mathbf{1}\{\mathbf{s}\in\mathcal{R}_{r}^{(t)}\}}.$ (33)
In contrast, GBT build classification trees sequentially, where each new tree is trained to optimize a loss function based on the residual errors of the previous trees (Friedman 2001). Let $R$ be the number of boosting iterations. Given covariates $\mathbf{s}$, for the $r$-th tree where $r\in[R]$, let $h_{r}(\mathbf{s})$ be the prediction and $\gamma_{cr}$ be the weight for class $c$. The predicted probability for class $c$ in a GBT classifier is obtained using the softmax function of the log-odds:
$\hat{p}_{c}(\mathbf{s})=\frac{e^{G_{c}(\mathbf{s})}}{\sum_{c^{\prime}=1}^{C}e^{G_{c^{\prime}}(\mathbf{s})}},$ (34)
with the log-odds defined as:
$G_{c}(\mathbf{s})=\sum_{r=1}^{R}\gamma_{cr}h_{r}(\mathbf{s}).$ (35)
15 Point Prediction ML Models
The regression models employed to generate point predictions are presented below.
  1. 1.
    Linear Regression (LinReg): LinReg solves the Ordinary Least Squares (OLS) problem.
  2. 2.
    Ridge Regression (Ridge): Ridge adds an $L_{2}$ norm regularization term to the coefficients obtained from OLS.
  3. 3.
    Least Absolute Shrinkage and Selection Operator Regression (Lasso): Lasso adds an $L_{1}$ norm regularization term to the coefficients obtained from OLS.
  4. 4.
    Random Forest (RF-Reg): RF-Reg operates similarly to the Random Forest classifier described in Section 14, but it builds a collection of regression trees instead.
  5. 5.
    Gradient Boosted Trees (CatBoost-Reg and XGBoost-Reg): Two versions of Gradient Boosted Trees (GBT) are implemented for regression. These GBT models are similar to the GBT classifiers described in Appendix 14, except they use regression trees. CatBoost-Reg is implemented using the CatBoost package, which builds symmetric trees and has native support for categorical features (Prokhorenkova et al. 2018). XGBoost-Reg is implemented using the XGBoost package (Chen and Guestrin 2016).
Table 5: Average Out‐of‐Sample MSE of Regression Models.
Model LinReg Ridge Lasso RF-Reg CatBoost-Reg XGBoost-Reg
MSE 2.324 2.324 3.149 1.791 1.668 1.666
16 Hyperparameters of the ML Models
Tables 6, 7, and 8 detail the hyperparameters and corresponding search spaces employed during the model selection and tuning processes for the ML algorithms considered in this study.
Table 6: Hyperparameters for Point Prediction Regression Models
Model Hyperparameter Range
Ridge $\alpha$ ($L_{2}$ regularization) $1e-4\leq\alpha\leq 10.0$
Lasso $\alpha$ ($L_{1}$ regularization) $1e-4\leq\alpha\leq 10.0$
RF-Reg n_estimators $100\leq\text{n\_estimators}\leq 1000$ (step: 100)
max_depth $3\leq\text{max\_depth}\leq 10$
min_samples_split $2\leq\text{min\_samples\_split}\leq 10$
min_samples_leaf $1\leq\text{min\_samples\_leaf}\leq 20$
CatBoost-Reg iterations $50\leq\text{iterations}\leq 300$
depth $4\leq\text{depth}\leq 10$
learning_rate $0.01\leq\text{learning\_rate}\leq 0.3$
l2_leaf_reg $1\leq\text{l2\_leaf\_reg}\leq 10$
bagging_temperature $0.0\leq\text{bagging\_temperature}\leq 1.0$
random_strength $0.0\leq\text{random\_strength}\leq 10.0$
border_count $1\leq\text{border\_count}\leq 255$
colsample_bylevel $0.5\leq\text{colsample\_bylevel}\leq 1.0$
XGBoost-Reg n_estimators $100\leq\text{n\_estimators}\leq 1000$ (step: 100)
max_depth $3\leq\text{max\_depth}\leq 10$
learning_rate $0.01\leq\text{learning\_rate}\leq 0.3$
subsample $0.5\leq\text{subsample}\leq 1.0$
colsample_bytree $0.5\leq\text{colsample\_bytree}\leq 1.0$
gamma $0\leq\text{gamma}\leq 5$
reg_alpha $0.0\leq\text{reg\_alpha}\leq 100.0$
reg_lambda $0.0\leq\text{reg\_lambda}\leq 100.0$
min_child_weight $1\leq\text{min\_child\_weight}\leq 10$
max_delta_step $0\leq\text{max\_delta\_step}\leq 10$
colsample_bylevel $0.5\leq\text{colsample\_bylevel}\leq 1.0$
colsample_bynode $0.5\leq\text{colsample\_bynode}\leq 1.0$
Table 7: Hyperparameters for Classification Models
Model Hyperparameter Range
MLR C (regularization) $1e-4\leq\text{C}\leq 1e2$
Classification Tree max_depth $3\leq\text{max\_depth}\leq 10$
min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 5)
RF-MC-CLF n_estimators $600\leq\text{n\_estimators}\leq 1000$ (step: 200)
max_depth $3\leq\text{max\_depth}\leq 10$
min_samples_split $2\leq\text{min\_samples\_split}\leq 10$
min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 5)
CatBoost-MC-CLF iterations $50\leq\text{iterations}\leq 300$
depth $4\leq\text{depth}\leq 10$
learning_rate $0.01\leq\text{learning\_rate}\leq 0.3$
l2_leaf_reg $1\leq\text{l2\_leaf\_reg}\leq 10$
bagging_temperature $0.0\leq\text{bagging\_temperature}\leq 1.0$
random_strength $0.0\leq\text{random\_strength}\leq 10.0$
border_count $1\leq\text{border\_count}\leq 255$
colsample_bylevel $0.5\leq\text{colsample\_bylevel}\leq 1.0$
Table 8: Hyperparameters for Quantile Regression Models
Model Hyperparameter Range
Regression Tree max_depth $3\leq\text{max\_depth}\leq 10$
min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 10)
QRF n_estimators $600\leq\text{n\_estimators}\leq 1000$ (step: 200)
max_depth $3\leq\text{max\_depth}\leq 10$
min_samples_split $2\leq\text{min\_samples\_split}\leq 10$
min_samples_leaf $10\leq\text{min\_samples\_leaf}\leq 30$ (step: 5)
17 Evaluation of the Simulator
The simulator is compared against a baseline that uses multinomial logistic regression (MLR), a common approach in the literature (Salari et al. 2022).
Table 9 reports the average CRPS across all carriers. The proposed calibrated MC-CLF simulator achieves a CRPS of 0.314, compared with 0.779 for the MLR baseline, indicating a substantial improvement in distributional accuracy at the individual‐instance level.
Figure 16 overlays each simulator’s joint predicted distribution on the true empirical distribution from the test set. The proposed calibrated MC-CLF simulator closely matches the empirical mass at zero deviation and reproduces both early‐ and late‐delivery tails, whereas the MLR baseline systematically overpredicts late deviations and underpredicts on‐time deliveries.
Figure 17 presents three carrier‐specific examples. In each case, the proposed calibrated MC-CLF simulator aligns tightly with the carrier’s empirical distribution across all deviation classes. By contrast, the MLR baseline exhibits clear misalignments—particularly in the majority classes and tails—underscoring the proposed simulator’s superior ability to capture carrier‐level heterogeneity.
Table 9: Average In-Sample CRPS Across All Carriers.
Model Calibrated MC-CLF Simulator MLR Simulator (Baseline)
CRPS 0.314 0.779
(a) Calibrated MC-CLF Simulator
(b) Baseline
Figure 16: Overall Joint Predicted Distribution of the Simulators vs. the True Empirical Distribution
(a) Example Carrier 1: Calibrated MC-CLF Simulator
(b) Example Carrier 1: Baseline
(c) Example Carrier 2: Calibrated MC-CLF Simulator
(d) Example Carrier 2: Baseline
(e) Example Carrier 3: Calibrated MC-CLF Simulator
(f) Example Carrier 3: Baseline
Figure 17: Selected Carrier Examples: Joint Predicted Distribution of the Simulators vs. the True Empirical Distribution
18 Additional Details on Computational Scalability
Table 10 provides a comprehensive comparison of objective values across different instance categories for the six proposed CSO methods. Table 11 summarizes the computation times for the baseline models and the proposed CSO-CLF methods.
Table 10: Realized Average ($\pm$ 95% CI margin of error) and Worst‐Case Objective Values Across Instance Categories.
Instance C-SAA-CLF C-SAA-QRF C-RO-B-CLF C-RO-B-QRF C-RO-D-CLF C-RO-D-QRF Avg Worst Avg Worst Avg Worst Avg Worst Avg Worst Avg Worst Small $0.821\pm 0.003$ 0.790 $0.848\pm 0.004$ 0.813 $0.851\pm 0.004$ 0.794 $0.871\pm 0.003$ 0.813 $0.936\pm 0.004$ 0.757 $0.920\pm 0.004$ 0.766 Medium $0.799\pm 0.008$ 0.932 $0.806\pm 0.009$ 0.925 $0.797\pm 0.009$ 0.834 $0.815\pm 0.009$ 0.821 $0.890\pm 0.010$ 0.822 $0.883\pm 0.009$ 0.825 Large $0.847\pm 0.012$ 1.032 $0.871\pm 0.017$ 1.055 $0.823\pm 0.014$ 0.982 $0.836\pm 0.016$ 0.981 $0.910\pm 0.017$ 0.972 $0.904\pm 0.014$ 0.973 Total $0.820\pm 0.003$ 0.823 $0.844\pm 0.003$ 0.842 $0.842\pm 0.003$ 0.811 $0.861\pm 0.003$ 0.824 $0.928\pm 0.003$ 0.779 $0.914\pm 0.003$ 0.787
Table 11: Computation Time Statistics (in seconds) Across Instance Categories.
Instance C-SAA-CLF C-RO-B-CLF C-RO-D-CLF C-Empirical-SAA Empirical-SAA PTO Greedy Avg Std Max Avg Std Max Avg Std Max Avg Std Max Avg Std Max Avg Std Max Avg Std Max Small 1.620 2.267 28.021 0.016 0.024 0.374 0.199 0.301 5.475 2.520 3.552 46.90 2.510 3.547 46.86 0.014 0.021 0.684 0.001 0.001 0.008 Medium 27.831 10.245 107.347 0.281 0.106 0.860 3.229 2.008 39.222 43.742 16.475 165.717 43.864 16.678 179.577 0.243 0.103 1.027 0.013 0.007 0.060 Large 59.266 22.590 182.401 0.698 0.322 2.856 7.683 7.056 111.188 91.190 35.100 278.677 92.225 35.421 295.212 0.629 0.329 2.584 0.052 0.034 0.317 Total 8.974 17.433 182.401 0.096 0.203 2.856 1.101 2.791 111.188 13.959 27.052 278.677 14.033 27.304 295.212 0.086 0.187 2.584 0.006 0.016 0.317
18.1 Sample Size Selection: Balancing Solution Quality and Computational Efficiency
Figures 18 and 19 illustrate the trade-off between average realized objective value and computation time for the two C-SAA methods.
Similarly, Figures 20 and 21 depict the trade-off between worst-case realized objective value and computation time for the two C-RO-D methods.
(a) Small Instances.
(b) Medium Instances.
(c) Large Instances.
Figure 18: Computation Time by C-SAA Sample Size.
(a) Small Instances.
(b) Medium Instances.
(c) Large Instances.
Figure 19: Average Realized Objective Value by C-SAA Sample Size.
(a) Small Instances.
(b) Medium Instances.
(c) Large Instances.
Figure 20: Computation Time by C-RO-D Sample Size.
(a) Small Instances.
(b) Medium Instances.
(c) Large Instances.
Figure 21: Worst-Case Realized Objective Value by C-RO-D Sample Size.